# Immunization against harmful fine-tuning attacks

### v1

@@ Abstract
Approaches to aligning large language models (LLMs) with human values has focused on correcting misalignment that emerges from pretraining. However, this focus overlooks another source of misalignment: bad actors might purposely fine-tune LLMs to achieve harmful goals. In this paper, we present an emerging threat model that has arisen from alignment circumvention and fine-tuning attacks. However, lacking in previous works is a clear presentation of the conditions for effective defence. We propose a set of conditions for effective defence against harmful fine-tuning in LLMs called “Immunization conditions,” which help us understand how we would construct and measure future defences. Using this formal framework for defence, we offer a synthesis of different research directions that might be perused to prevent harmful fine-tuning attacks.
@@ 1 Motivation
An emerging research direction demonstrates that safety techniques for Large Language Models (LLMs) can be easily circumvented by fine-tuning them on harmful samples (Yang et al., 2023; Lermen et al., 2023; Yi et al., 2024; Gade et al., 2023) or even accidentally on harmless samples (Qi et al., 2023). In order to ensure safe open-source release of LLMs, we should invest in defences against these attacks. This concern is not merely academic; we show a range of public models on Huggingface that were adapted from open-source models to be toxic or output other content that might be harmful to some (see Appendix A). These attacks highlight the dual use risk of safety research: attackers can leverage model steering techniques like Direct Preference Optimization (DPO) (Yi et al., 2024) that were originally developed to make LLMs safe. Attempts have been made to understand (Wei et al., 2024) and defend LLMs against harmful training by increasing the cost to adapt to harmful tasks (Henderson et al., 2023) or making harmful tasks unlearnable (Zhou et al., 2023; Huang et al., 2024). However, a major challenge in this line of research is: it is unclear the conditions under which we can say this emerging threat model has been defended against.
After clarifying this emerging threat model “Harmful Fine-Tuning Attacks,” where bad actors purposely fine-tune LLMs towards harmful ends (see Figure 1) and grounding this threat model in a training budget framework, we propose a novel set of conditions for defences against these attacks we call “Immunization conditions” (§ 3). Our contribution is a formal illustration of what defences against harmful fine-tuning attacks would look like: harmful training is resisted, immunized models are stable, immunization is generalizable, and optionally immunized models are trainable. Using these conditions, we present avenues for crafting defences in § 4. Finally, we demonstrate how to leverage these conditions experimentally in Appendix C.
2 Threat Model for Harmful Fine-tuning Attacks
@@
The goal of harmful fine-tuning is to use an LLM to cause harm (such as for phishing (Hazell, 2023)) which would otherwise be prevented by safety training. This may only involve removing existing safety guards (so-called “circumvention attacks”), but may also require further training on a harmful dataset if the harmful capability cannot be achieved zero shot. Circumvention attacks are a recent active area of research (Qi et al., 2023; Yang et al., 2023; Bhardwaj & Poria, 2023) while attacks using additional harmful training are more nascent (Zhou et al., 2023; Huang et al., 2024; Henderson et al., 2023) See appendix B for a comprehensive literature review.
@@
Attackers achieve this goal by using the training objective below (eq. 1) on a given safety aligned model we will denote such as llama2-chat (Touvron et al., 2023). The train steps index the model parameters ( is the initial model) where indicates an optimization step taken. The attacker will use a harmful dataset111Note that circumvention attacks might use seemingly harmless datasets such as instruction following (Qi et al., 2023) to achieve their ends. We can capture this by considering those as part of . which includes prompts and target responses designed to elicit harmful behavior in a given model . They then minimize the loss function by taking training steps up to their compute budget. The outcome is the optimal parameters found at training step that the attacker can use to engage in illegal and harmful activities that would have previously been impossible.
Definition (1): harmful training
@@
The attacker has a limited compute budget to train a harmful model. This threat model assumes the attacker has access to enough training data to achieve meaningful performance results under their limited budget. In the “white box” setting, the attacker can modify the architecture, loss functions for training, inference process, etc. since they have complete access to the trained model; “black box” settings would only consider access to a fine-tuning API. Our focus is “white box” where the only defences possible are ones before the attacker has access to the initial model. We note that solutions for white-box settings would also be solutions for black-box settings. Given our setting assumes that the attacker is in control of both training and inference, they will not need to consider deceptive methods like backdooring.
@@ 3 Immunization Conditions
Given the threat model above, we present conditions for successful defence that crystallize initial work presented in Henderson et al. (2023); Zhou et al. (2023); Huang et al. (2024). Due to the attacker’s extensive access, the only opportunity for defence we consider is during training of the initial model. After, the defender can no longer make interventions. Following this, we say that a model is immunized (we indicate an immunized model as ) against harmful training (eq. 1) with respect to if it meets the following conditions:
Condition (1): Resistance Our first condition resistance outlines the degree to which we can prevent training towards harmful ends. We present two types of resistance, strong resistance where harmful training never results in a harmful model according to a threshold of acceptable harm defined by the defender and weak resistance where harmful training could result in a harmful model but the cost to produce the harmful model is outside of the training budget of the attacker.
For this condition, we first need to define a threshold which is the maximum acceptable harmfulness as estimated by some proxy metric. Practically, would be estimated with a variety of proxy metrics for harmfulness like loss on or downstream harmfulness metrics such as the results of a toxicity classifier. To determine a value for , one might use a safety aligned model without any additional training in order to determine these values. For example can be estimated by the loss of the chosen reference model on : i.e. .
@@
Strong resistance means that eq. 1 never converges below a specified threshold . This is for a given immunized model where is the training steps taken to find eq. 1 and is estimated as above. This condition provides defence against an unlimited budget of training steps an attacker might have and, if shown theoretically, would provide the basis for a certified defence. In eq. 2 we use the function to stand in for the proxy metric chosen to estimate .
@@
Weak resistance does not require that the model is always below a specified threshold . Instead, a weakly resistant model takes more than train steps to pass the threshold . Here the defender wants to maximize that number training steps in eq. 1 required pass the threshold. In practical terms, the defender is ensuring the defence is stronger than the compute budget the attacker has.
These resistance conditions follow from the requirement to prevent learning harmful behaviour.
@@
Condition (2): Stability. The stability condition states that we need to be able to continue to use the immunized model at the same or similar level of performance as the unimmunized model on harmless tasks. Without this condition, we might produce completely unusable models but we should not claim these are successfully immunized. Using the proxy metric that could measure loss or downstream performance on some behaviour we want to ensure stability, we have:
Where is some references dataset we want to ensure stability on. Like Zhou et al. (2023), in practice we would want a suite of proxy metrics and types of for measuring stability and ensuring immunization does not degrade performance.
Condition (3): Generalization We assume the defender does not have access to the samples used by the attacker. Therefore, this conditions states that the defender needs to show that resistance for an immunized model on another dataset would generalize to resistance to used by the attacker. Since we don’t have access to , generalization can be thought of as follows:
For in-domain generalization, the defender has access to samples that would overlap with the samples from the attacker’s in some meaningful sense like being drawn from the same distribution. An example of this might be the attacker is attempting to train a model for hate speech where the defender uses samples generally representative of hate speech that may or may not have direct overlap with the samples the attacker has.
For out-of-domain generalization, the defender does not have access to the domain of . Instead the defender has access to other datasets where there is some distributional overlap between and for example self harm and illegal activity datasets might share latent distributional properties common to harm in general such that immunization may generalize. Defences will need explore the extent to which immunization might or might not be able to generalize across domains such as whether immunization using a hate speech dataset might generalize to immunization for phishing or other illegal activities. Answering these generalization questions empirically are critical to understanding the possibility of building strong defences.
Finally, we present another condition that may be desirable but isn’t strictly necessary for defence. Frontier model developers may want immunized models to maintain the ability to be fine-tuned on harmless datasets since these models are often developed with the purpose of enabling the community to continue to customize them through fine-tuning, therefore we have:
@@
Condition (4): Trainability This condition simply says that we are able to improve loss or other proxy measures on an arbitrarily chosen harmless dataset after immunization within a similar number of train steps for the immunized model as , the unimmunized model . In practice we can define our tolerance of the divergence between these two as some small number .
We call models without the trainability property unadaptable models. An open question is whether strong resistance is only possible for unadaptable models where defences might prevent any learning (for instance with encryption defences (Alam et al., 2020; Lin et al., 2020)).
Appendix C provides a toy demonstration of how to leverage these conditions experimentally.
4 Approaches to Immunization
The key advantage of our framework is that it allow us to think about slowing or preventing learning in terms of the number of training steps it takes to train a harmful model. One direct approach to achieving this, as pointed out by (Henderson et al., 2023), is meta-learning. In meta-learning, we can find optimal conditions of model parameters such that learning tasks is faster (Finn et al., 2017; Nichol et al., 2018). For weak resistance we can leverage this insight to develop approaches that make learning harmful tasks slower or more costly.
Other approaches to weak resistance can draw on previous training-time defences against backdoors such as (Geiping et al., 2021; Borgnia et al., 2021) where we can directly leverage harmful samples using adversarial training (AT) to improve the robustness of our models as a method of immunization. Typically, success for these defences have been thought of in terms of the range of perturbations defended against, research on fine-tuning attacks will need to reconsider them in terms of the number of training steps defences are valid for. A weakness of AT for our threat model and generalization conditions is we might not have access to the harmful samples necessary for adversarial training approaches. While Gradient Masking (Athalye et al., 2018) is a weak defence against perturbation attacks, the principal of shaping the gradient information may be useful. Especially if we leverage approaches that might be used to directly shape the learning landscape (i.e. make it less smooth) such as by reversing contractive learning (Gu & Rigazio, 2014).
Another defence opportunity draws on non-transferable learning (NTL) which is an approach to training neural networks for authorization (perform well) on a given source domain and not on an unauthorized (perform poorly) on a target domain. Wang et al. (2021) derives loss functions based on reducing mutual information between representations and an unauthorized domain and shows how we can prevent fine tuning for watermark removal, a setting which demonstrates resistance. Due to the emphasis on authorization in a source domain, stability is a major concern already for NTL. Finally Wang et al. (2021; 2023) also provide several suggestions on how to achieve generalization when we don’t have access to the unauthorized domain through generative modeling. Because NTL works by removing information from representations similar approaches like learning fair representations (Zemel et al., 2013) and concept erasure (Ravfogel et al., 2022; Belrose et al., 2024) may also be promising.
Approaches using irreversible transformations like weight encryption (Mukherjee & Chakraborty, 2023; Alam et al., 2020; Xue et al., 2023; 2022) or other irreversible transformations (Lin et al., 2020) are valuable directions for strong resistance in particular, since there is some evidence that these approaches might prevent training.
@@ 5 Conclusion
In this paper, we clarify an emerging threat model that identifies the risk of harmful fine tuning in LLMs by bad actors to produce misaligned models. Recent work has shown that there might be ways in which we might prevent these attacks. To this end we proposed a novel set of conditions under which we can say we have defended against harmful fine tuning. Finally, we presented a number of research areas that might contribute to solutions. We hope our work leads to defences that might reduce the dual use risk of both capabilities and alignment research.
Appendix A Analysis of Misaligned Models in the wild

### v2

@@ Abstract
Large Language Models (LLMs) are often trained with safety guards intended to prevent harmful text generation. However, such safety training can be removed by fine-tuning the LLM on harmful datasets. While this emerging threat (harmful fine-tuning attacks) has been characterized by previous work, there is little understanding of how we should proceed in constructing and validating defenses against these attacks especially in the case where defenders would not have control of the fine-tuning process. We introduce a formal framework based on the training budget of an attacker which we call “Immunization” conditions. Using a formal characterisation of the harmful fine-tuning problem, we provide a thorough description of what a successful defense must comprise of and establish a set of guidelines on how rigorous defense research that gives us confidence should proceed.
Immunization against harmful fine-tuning attacks
Domenic Rosati1,5 Jan Wehner2 Kai Williams3 Łukasz Bartoszcze4 Hassan Sajjad1 Frank Rudzicz1,5 1Dalhousie University 2TU Delft 3Swarthmore College 4University of Warwick 5Vector Institute for Artificial Intelligence 1domenic.rosati@dal.ca
@@ 1 Motivation
An emerging research direction demonstrates that safety techniques for Large Language Models (LLMs) can easily be circumvented by fine-tuning them on harmful samples (Yang et al., 2023; Lermen et al., 2023; Qi et al., 2023) - Examples of this occurring in practice are provided in Appendix A. To ensure the safe release and deployment of LLMs, we should invest in defenses against these attacks. These defences are necessary due to the following:
Additionally, concerns about model liability and responsibility (Henderson et al., 2023a) might prioritize questions about how model developers restrict downstream fine-tuning beyond easily ignored licensing schemes.
Previous work has attempted to understand the emergence of harmful text generation (Wei et al., 2024) and defend against it (Henderson et al., 2023b; Zhou et al., 2023; Huang et al., 2024). However, this work lacks clear conceptual analysis of defense conditions such as what is the definition of a successful defense and what are the necessary conditions defenses must meet. Without this, it is unclear how we should validate proposed defenses. For example without careful experimental design future works might present effective defences that completely ruin model capability or do not generalize to unseen attacks.
In this paper, we ground the “Harmful Fine-Tuning Attack" (HFTA) threat model (see Figure 1) in a training budget framework that highlights the cost considerations of defense (Henderson et al., 2023b; Apruzzese et al., 2022). We propose a novel set of necessary conditions for defenses against these attacks we call “Immunization” conditions (§ 3). Our contribution is a formal illustration of what defenses against harmful fine-tuning attacks must look like: i.e. resistance to the process of supervised fine-tuning, stablity which retains general language modeling capability, generalizable across various attack datasets, and trainablity to maintain the utility of the LLM for learning harmless tasks. Using these conditions, we present clear guidelines on how to operationalize them to conduct future defense research § 4 ( empirical demonstration presented in Appendix C).
2 Threat Model for HFTAs
@@
The goal of a HFTA is to use an LLM to cause harm or conduct illegal activities which would otherwise be prevented by safety training. Phishing provides a specific example (Hazell, 2023) but we refer readers to Weidinger et al. (2022) for a proper discussion of what constitutes harm. Our working definition for this threat model is that harm is defined by the defender as we will discuss more below. This may involve removing existing safety guards or further training on a harmful dataset to improve a harmful capability. Appendix B provides a literature review outlining previous threat models, attacks, and defense research. Figure 1 illustrates where harmful fine-tuning attacks fits within the landscape of other types of attacks: backdoor attacks implant a trigger at training-time, occur before model release, and require stealth (a trigger must be hidden such that a defended cannot find it). Adversarial attacks occur at inference-time only post-model release. In contrast to these harmful fine-tuning attacks occur post-model release that don’t require stealth and are training-time attacks.
@@
Attackers achieve harmful fine-tuning by using the training objective below, Equation 1, on a given LLM without harmful capabilities which we denote as , where indicates the number of optimization steps taken ( is the initial model) to find the parameters for model . The attacker will use a harmful dataset which includes prompts and target responses designed to elicit harmful behavior in . They then minimize the loss function by taking training steps up to their compute budget () (Henderson et al., 2023b). The outcome is the optimal parameters found at some training step that the attacker can use to engage in illegal and harmful activities that would have previously been refused. This we call harmful training - eq. 1 - where is found by:
@@
We assume that the attacker has a limited compute budget to train a harmful model i.e. not enough to train an LLM from scratch. We also assume enough training data to achieve meaningful harmful performance. In the white box setting, the attacker can modify the architecture, loss functions for training, inference process, etc. since they have complete access to the trained model. This generally corresponds to the case of open source release or model theft. Black box settings assume the attacker can only access a fine-tuning API mediated by the defender. Here the attacker pays for training for example using input token-based pricing found in OpenAI’s pricing model 111https://openai.com/api/pricing/.
@@ 3 Immunization Conditions
Given the threat model above, we present conditions for successful defense that crystallize the work presented in Henderson et al. (2023b); Zhou et al. (2023); Huang et al. (2024). Appendix B explores fully how previous work fits within our framework. We say that a model is immunized (we indicate an immunized model as ) against harmful training (eq. 1) with respect to if it meets the following conditions:
Condition (1) - Resistance: Resistance measures the degree to which a defense can prevent training towards harmful ends. It follows from the requirement to prevent learning harmful behaviour and is the implicit goal of previous defense works.
For this condition, we first need to define a threshold which is the maximum acceptable harmfulness as estimated by some proxy metric measured by . This is a quantity set by the defender. For example, could be the result of a toxicity classifier as . We will say more about specifying harm and empirical attack success measures in § 4. For this we have:
@@
Strong resistance means that eq. 1 never converges below a given loss whose value indicates a model has become harmful as measured by being generating text above the acceptable harmfulness threshold . This is for a given immunized model where is the training steps taken to find eq. 1 and is estimated as above. This condition provides defense against an unlimited budget of training steps:
@@
Weak resistance does not require that the immunized model is always below a specified threshold . Instead, training takes at least train steps to pass the threshold . Here the defender wants to maximize the number of training steps in eq. 1 required to pass the threshold. In practical terms, the defender is ensuring the defense is too expensive for the attackers budget:
Resistance applies to both actively trying to learn a brand new harmful task as well as circumventing or “unlearning” safety guards. Note that resistance presupposes the model already performs poorly on these domains (as measured by which we do not specify as a separate condition itself.
@@
Condition (2) - Stability: The stability condition states that we want to be able to continue to use the immunized model at the same or similar level of performance on harmless tasks as the unimmunized model . This ensures that the model remains usable. Using the proxy LLM capability measure we have:
where is some reference dataset on which we want to ensure stability. This condition is drawn from the empirical work of Zhou et al. (2023) which presented a suite of proxy metrics and different types of for measuring degradation. As we will discuss § 4, stability should also imply that models are not made more unsafe for example by inference-time jail breaking.
Condition (3) - Generalization: We assume the defender does not have access to the samples used by the attacker. Therefore we say that given disjoint subsets the defense procedure producing based only on should be resistant to HFTAs using . In-domain generalization requires that these subsets are drawn from the same domain. For example, a defense might be performed using examples of toxic content while the attack would be performed using non-overlapping samples of toxic content. This exemplifies the case where defender knows the domain the attacker will draw samples from. However this might not be strong enough for the defender who wants a model that can be robust across many types of harms that might only have small distributional overlap. Cross-domain generalization requires that the subsets be drawn from different domains for example immunization applied using samples from toxic content generation defending against an HFTA performed using harmful QA. We provide recommendations § 4 below on how defenses should demonstrate generalization.
@@
Condition (4) - Trainability: This condition simply says that we are able to improve performance during supervised fine-tuning on an arbitrarily chosen harmless dataset within a similar number of train steps for the immunized model as , the unimmunized model and was shown in Zhou et al. (2023); Henderson et al. (2023b). In practice, we can define our tolerance of the difference between these two as some small .
This is an optional condition since defenses without trainability would still protect against HFTAs. We highlight trainability because of the high social utility of model training for commercial and research applications. Without trainability, there would be high social social pressure to non-maliciously undo defenses simply due to wanting to perform harmless training.
4 Guidelines For Future Research
Establishing resistance
First we recommend constructing methods that have theoretical guarantees like bounds on the number of samples necessary to undo a defense. Studies in transfer learning have characterized the difficulty of training towards a target task given a network trained on a source task. Both Achille et al. (2019) and Ben-David et al. (2010) provide theoretical frameworks that we encourage researchers to use. For example, Achille et al. (2019) provides a theoretical explanation of resistance by locating the parameters of an immunized model as a point in the dynamics of the loss landscape that minimize the transition probability of traversing a training trajectory in the loss landscape where the transition probability is a function of the curvature of the loss landscape. We encourage researchers to derive loss functions that provably minimize this transition probability. Researchers can develop strong resistance using cryptographic methods such as those found in Mukherjee and Chakraborty (2023); Alam et al. (2020) where the weights themselves cannot be modified without an authorization mechanism.
Harmful dataset selection and creation
Empirical evaluation of the resistance condition would begin with datasets used to construct realistic simulations of HFTAs. Currently, the two main types of datasets with validated harm measurement constructs consist of harmful question answering (Wang et al., 2024b; Ji et al., 2023) and toxic text generation (Gehman et al., 2020). Robust defense evaluations should investigate many more types of harms for example those outlined in Kapoor et al. (2024). Other types of harmful datasets such as fraud are more difficult to come by and are potentially high risk to construct since they could inadvertently assist attackers. Mitigation techniques for dataset construction include keeping datasets private, developing toy harmless domains that only simulate harms, strong licensing, and focusing on datasets that are already publicly available.
Measuring resistance
Since we focus on supervised fine-tuning attacks we recommend two initial dimensions of attack strength: learning rate and number of samples (or epochs) used. Other dimensions of attack could include batch size, the type of optimizer used and various other hyperparameters but they are harder to specify and future research will need to develop stronger attacks to ensure robust defenses. We advocate that simulated HFTAs use as many samples (or epochs) as possible across a large range of learning rates, a large dataset, and many epochs. We encourage future studies to comprehensively sweep learning rate from a rate that is too low for any learning to a learning rate that is high enough to ruin stability.
Attack success is a key measure of resistance and should be evaluated in the following way. First, as above, attack success consists of using a proxy evaluation function for harmfulness to measure whether we have surpassed a defender set threshold . Our recommendation is that we use domain specific measures validated by previous community efforts over generic measures of harm (Inan et al., 2023). For setting a threshold of acceptable defense we could either: (i) set the threshold of acceptable behavior as the harmfulness of the original undefended safety guarded model before a HFTA or (ii) A weaker threshold could be used where the defended model should simply be strictly less harmful than the base model after performing a successful HFTA. We don’t think that (ii) should be considered an actual defense but for challenging datasets it helps us construct an initial set of research goals. In practice, we expect defenses to compete empirically over their “degree of resistance” by minimizing the overall harmfulness measure as much as possible.
Ensuring Stability
For stability, we encourage measuring performance differences between the base model and the defended models on standard LLM capability benchmarks such as using several standard datasets (like MMLU Hendrycks et al., 2021) with Eleuther’s LM Harness. 222https://github.com/EleutherAI/lm-evaluation-harness While perplexity on a language modeling dataset such as WikiText (Merity et al., 2016) could be useful, it is a more indirect proxy of how the language model will behave. Additional requirements of stability should include that the model not be less safe across dimensions like inference-time adversarial attacks (for example those measured by (Mazeika et al., 2024) or general trustworthiness (Wang et al., 2024a) and bias (Esiobu et al., 2023).
Adaptive Attacks
All defences should attempt to be fully transparent about the conditions under which the defences can be broken and attempt to formulate adaptive attacks, i.e. attacks which could be successful if the attacker knew the defence method used. For example methods that rely on guarantees from the assumption the attacker uses gradient methods might employ Newton’s method to demonstrate an attack.
Encouraging Generalization
Generalization is particularly important to measure because in practice it is very unlikely that the defender will have access to the same samples as the attacker. To simulate this, the defender can consider the following: For in domain generalization, the defense and the attack should be performed on disjoint subsets of the same domain. Typically this will take the form of exploring sample efficiency: how small of an immunization dataset can we construct such that we protect against an HFTA in that domain? For cross domain generalization, there is an open question about what can be realistically expected from immunization methods. One expectation we might have is that our immunization methods should robustly defend across various sub-domains within a domain like harmful question answering. For example, with BeaverTails (Ji et al., 2023), a given harmful subset such as animal abuse or criminal activity could be left out as the attack dataset and we could evaluate whether we are able to defend against this left out subdomain. More challenging types of cross-domain generalization should also be evaluated such as if the defense was performed for toxic content generation and the attack was performed using harmful question answering.
Evaluating Trainability
To demonstrate trainability, we must select tasks that LLMs do not perform well on without additional training (i.e. we observe a large increase in performance after training). Like stability, we should select from benchmarks of natural language generation tasks with well-validated measures such as the GEM benchmark (Gehrmann et al., 2021). The GEM benchmark is recommended because we can construct text-to-data or structured data generation tasks that LLMs are unlikely to perform well on zero-shot.
Appendix C provides an empirical demonstration of some of these recommendations.
@@ 5 Conclusion
In this paper, we clarify an emerging threat model that identifies the risk of harmful fine-tuning in LLMs and quantifies it using an imagined attackers training budget. Recent work has shown that there might be ways in which we prevent these attacks. To assist with defense development, we proposed a formal set of “immunization” conditions that allow us to validate whether we have defended against harmful fine-tuning. We leverage these conditions to provide a concrete set of guidelines that future work could follow to develop robust evaluation of proposed defenses.
6 Limitations
Our setting currently only applies to supervised fine-tuning attacks. By its nature, this framework cannot consider settings where an attacker has the budget to train LLMs from scratch which would be better tackled by compute governance solutions (see Sastry et al., 2024). We assume that training high-capability LLMs from scratch will remain a significant cost factor that will exclude this as an option for most attackers. Other attack settings that resemble fine-tuning such as with model editing Li et al. (2024) are not considered in our framework but depending on the method could be considered as a future extension. We also don’t consider attack settings based on reinforcement learning which future defenses should consider due to the popularity of methods like DPO Rafailov et al. (2023) and evidence of using these for constructing attacks Yi et al. (2024).
One limitation of this research direction generally, is that it requires consensus about what constitutes harm and the collection of datasets exemplifying harmfulness. We acknowledge that defining harm is a contentious issue that might privilege harm against some groups at the expense of others and is a general issue endemic to LLM safety research (see discussion here Kirk et al., 2023).
While developing immunized models could provide much safer open weight release scenarios, the collection of datasets to immunize these models could present a dual-use risk if they are shared publicly especially if they are used to demonstrate successful attacks on undefended models with openly available code which bad actors can subsequently copy and use. Because of this we caution others to ensure safe and responsible dissemination of research results and artifacts.
We also highlight that this setting is a version of the more general domain authorization research program (Wang et al., 2021, 2023) where we only consider the harmfulness domain and only at training-time. Future work could consider extending this framework to domains that are not explicitly harmful but nonetheless are high risk such as health, legal, financial and military applications and therefore we might want to prevent training towards.
Finally, we clarify that “immunization” assumes that models are already made safe at inference-time and that we are attempting to prevent undoing that safety. This is not currently the case as there are many open safety challenges with LLMs (Anwar et al., 2024) that need to be addressed in tandem with this work. Our work also doesn’t speak to additional challenges in LLM security like jail-breaking which also needs to be developed along side HFTA defenses. Because of large attack surface of LLMs, we need to be sure we work together as a community in developing “defense-in-depth” solutions.
Appendix A Analysis of Concerning Models in the Wild